Question 12 (1 point)

Suppose that a researcher studying the weight of female college athletes wants to predict the weights based on height, measured in inches, the percentage of body fat of an athlete, and age. The researcher calculates the regression equation as (weight) = 4.73*(height) + 1.45*(percent body fat) - 0.712*(age) - 84.809. If a female athlete is 64 inches tall, has a 19 percentage of body fat, is 19 years old, and has a weight of 235.297, the residual is 3.364. Choose the correct interpretation of the residual.

Question 12 options:

1)

The weight of the athlete is 3.364 pounds larger than what we would expect.

2)

The height of the athlete is 3.364 inches less than what we would expect.

3)

The weight of the athlete is 3.364 pounds less than what we would expect.

4)

The height of the athlete is 3.364 inches larger than what we would expect.

5)

The weight of the athlete is 235.297 pounds larger than what we would expect.
Answer #1

(weight) = 4.73*(height) + 1.45*(percent body fat) - 0.712*(age) - 84.809.

= 4.73*64 + 1.45*19 - 0.712*19 - 84.809.
= 231.933

a weight of 235.297,

Residual = 235.297 - 231.933

the residual is 3.364.


The weight of the athlete is 3.364 pounds larger than what we would expect.